Compute the future value of a $180 cash flow for the same combinations of rates and times: (Do not round intermediate calculations. Round your answers to 2 decimal places.)

     Future Value
  a. r = 12%, t = 8 years $       
  b. r = 12%, t = 16 years       
  c. r = 6%, t = 8 years       
  d. r = 6%, t = 16 years       
Answer #1

FV = PV(1 + r)t

a. FV = $180 × (1.12)8 = $445.67

b. FV = $180 × (1.12)16 = $1,103.47

c. FV = $180 × (1.06)8 = $286.89

d. FV = $180 × (1.06)16 = $457.26